OpenStudy (anonymous):

Can someone please check my work to make sure I understand. "Solve for variable specified" 1. -14n+q=rt-4n (solve for n) I combine n and since both are (-) I add. -18n + q = rt then I subtract q from each side. -18n+q-q=rt-q. So, -18n=rt-q. I then I add 18 to both sides and the answer is: n = rt-q+18 .... Yes?

OpenStudy (anonymous):

rosimary, this is an equation meaning the left and the right expressions are equal. You need to +4n to both sides...etc :)

OpenStudy (anonymous):

oh is rt, a square root sign?

OpenStudy (anonymous):

-10n+q=rt q-rt=10n q-rt/10= n

OpenStudy (anonymous):

q-rt is all over 10

OpenStudy (anonymous):

no there is no square root. it's Algebra 1. The teacher stated on the worksheet the following problems. "Solve for Variable Given" (Hint: Get variable on one side 1st) 1. -14n + q = rt - 4n Solve for "n" 2. 18t + 11v = w - 13t Solve for "t" 3. 8k + 9 = 7k + 6 "Solve Problem"

OpenStudy (anonymous):

rosimary use ur basic principles.... add, subtract, multiply, divide. get n to one side and everything to the other. dont make it harder that it has to b

OpenStudy (anonymous):

Got it. So did I get it right the 1st time? Problem 1. -18n + q = rt .... Then -18n + q-q = rt - q .... Then -18n= rt -q ... Then we add 18 to both sides which lets "n" stand alone and ANSWER: n = rt -q + 18

OpenStudy (anonymous):

q-rt/10= n q-rt is all over 10

OpenStudy (anonymous):

the work is above a little

OpenStudy (anonymous):

1. (add 4n to both sides) -10n + q = rt (minus q from both sides) -10n = -q + rt (divide by -10) \[\Large n = \frac{q}{10} -\frac{rt}{10}\] or \[\Large n = \frac{q-rt}{10}\] ^^as habibi said :)

OpenStudy (anonymous):

Do you understand rosimary? I think you thought the problem was -14n + q + rt - 4n, where you would add like expressions, but it was -14n + q = rt -4n, so it's an equation and you need to do the same thing to each side.

OpenStudy (anonymous):

Hi Josh. I think I got it. I guess I'm confused about the add like expressions. Because it is -14n + -4n wouldn't that be -18n since they are both negatives we add them? Thank you for showing the work step by step because that's how the teacher wants it especially on the test.

OpenStudy (anonymous):

rosimary, if u hav 2 negative signs then u add... say -14 and -4.... it is then -10 if u hav a neg and a posative then u hav to subtract and 2 pos u add

OpenStudy (anonymous):

i remember it by saying the 2 negative signs if put together looks like an addition sign

OpenStudy (anonymous):

Hi Rosimary :) Your absolutely right, two negatives do equal a positive! it's really great that you want to understand the process. If there wasn't an equals sign and it was, for example: 10n + 4n -3n , then you would add them together and it would equal 11n, but because we are using an equals sign in the middle, it means it's an equation, which is a fancy way of saying what is on both sides is equal. for example 4n + 1 = 3n + 4 , so you can't add the 3n to the 4n because we are saying that 4*(whatever n is) + 1 = the same number that 3*(whatever n is) + 4, so n is the same number on both sides and we have to find n. So, for my example if we subtract 3n from the right, we also have to subtract 3n from the left, so that they will still be equal. 4n + 1 = 3n + 4 -3n -3n (these 3n's will cancel out on the right cause 3n - 3n is 0) (4n - 3n = 1n) 1n + 1 = 4 - 1 -1 (we want the n by itself or isolated on one side so we can see what it is) (1 - 1 = 0, so they cancel out) 1n = 3 (/1) (/1) (divide both sides by 1 so we get the n by itself on the right) n = 3/1 (3 divided by 1 is just .... 3! ) so the answer would be n = 3, then you can put the three back into both sides of the first equation to see if it's true: 4*(3) + 1 = 3*(3) + 4 (replace all the n's with a 3 - because that's what we worked out n to be equal to) 12 + 1 = 9 + 4 or 13 = 13 (right!! it works. They are equal) I hope you understand :)

OpenStudy (anonymous):

I get it!!!! Yes we are learning equations. I spent the last 2 weeks hospitalized and missed a lot of work. Now I'm trying to catch up and I am extremely overwhelmed. I feel dumb and I found this site by the grace of God. Thank you. Let's see if I can I apply what you explained and you can check my work if you don't mind....

OpenStudy (anonymous):

That's really great that you get it! :) Sorry to hear that you were in the hospital, I hope your all better now. This site is really good, and also www.khanacademy.com is a really good site, he has a whole playlist on equations and he explains really easily. You shouldn't feel dumb, because you want to learn, and lots of people don't. They just want the answers. Your very welcome. Yes that would be fine :) i'll check your work.

OpenStudy (anonymous):

-14n + q = rt - 4n solve for n. so add 4 to each side. -10n + q = rt then we divide -10 to both sides. n + q = rt/10 then we subtract q to both sides. so answer is: \[n = \left(\begin{matrix}rt-q \\ 10\end{matrix}\right)\] .... Yes?

OpenStudy (anonymous):

Ok .... Just noticed your answer from earlier so I get it now.

OpenStudy (anonymous):

Yes! you're doing really well! one thing i forgot to mention is that it's better to wait until you have the -10n by itself on one side before we start dividing. so after you added 4 to both sides, it would be better to subtract the q from the left before you start dividing by what's on your target variable (n).

OpenStudy (anonymous):

josh i dont think it does..... the Q itself is the negative part so wouldnt it b -q+rt?

OpenStudy (anonymous):

sorry it's not. I'm tired :) q - rt is not the same as rt - q

OpenStudy (anonymous):

#2. 18t + 11v = w - 13t. We add 13 to each side. so turns into 31t + 11v = w. Then we subtract 11v to each side. turns into 31t = w - 11v. Then divide by 31 and the answer is \[t = \left(\begin{matrix}w-11v \\ 31\end{matrix}\right)\]

OpenStudy (anonymous):

yup you're right habibi, my mistake. I was thinking of the commutative law of multiplication/addition

OpenStudy (anonymous):

i see where ur coming from there. @rosimary i think thats right im tired myself too

OpenStudy (anonymous):

yup your right rosimary! well done girl :)

OpenStudy (anonymous):

me too. I have to take part 1 of my test in 5 hours. I'm just going to tell my teacher how overwhelmed I feel and hope she gives me more time to get all of this. I'm 30 going to back to High School to make up for my past mistakes. I didn't realize how challenging it would be. Thank you both so very much!

OpenStudy (anonymous):

well you're definitely on the right track, It just takes some time to sink in. You should check out khanacademy for sure though, he has taught me so much. Let us know if we can help you again and good luck with the test. I know what it's like. I was pretty messed up when I was younger, but now I'm teaching myself like you are :) good night guys

OpenStudy (anonymous):

I have a couple more math problems that will be on the test. "Clear Fractions First" 1/4 - 2/3y = 3/4 - 1/3y .... is this the same principle?

OpenStudy (anonymous):

well im in geometry myself so give me a heads up when u need help :)

OpenStudy (anonymous):

okie rosimary on this one u hav got to do like fractions.

OpenStudy (anonymous):

I'll leave you in Habibi's capable hands, I gotta get to bed :D night guys

OpenStudy (anonymous):

so 1/4 and 3/4 and 2/3 and 1/3 preferably id put everything to one side and set it equal to 0

OpenStudy (anonymous):

night josh

OpenStudy (anonymous):

Good night Josh.

OpenStudy (anonymous):

heres what i would do

OpenStudy (anonymous):

1/4 - 2/3y = 3/4 - 1/3y -2/4-1/3y=0 -2/4=1/3y -6/12=4/12y -10/12=y

OpenStudy (anonymous):

u need like fractions so the denominator has to b the same. i made it 12 that came to my head fastest

OpenStudy (anonymous):

ok let me look at it and see if I understand

OpenStudy (anonymous):

Habibi I know you are tired and I'm sorry to bother you but I don't understand. What did we subtract or add to both sides?

OpenStudy (anonymous):

its fine um i moved everything t the one side... and combined the like fractions. then moved the denominator of 3 back with y then i made the denominators like terms

OpenStudy (anonymous):

1st i subtracted 3/4 from both sides which left me with -2/4-2/3y=-1/3y

OpenStudy (anonymous):

then added 1/3y to both sides and made it -2/4-1/3y=0

OpenStudy (anonymous):

then added 1/3y back to the other sided so -2/4=1/3y

OpenStudy (anonymous):

u got there?

OpenStudy (anonymous):

@rosimary u there still?

OpenStudy (anonymous):

uh oh. poor rosimary

OpenStudy (anonymous):

im still here .... sorry

OpenStudy (anonymous):

okie did u want me to keep going?

OpenStudy (anonymous):

i can stay up a lil longer

OpenStudy (anonymous):

I'm extremely tired. I don't get it. I'm going to put this problem to the side and get at least 4 hrs of sleep before the test. I will tell my teacher to let me study for the 2nd part of the test one more day. I talk to you tomorrow. You go to sleep. Thank you and have a blessed night.

OpenStudy (anonymous):

i shall. and dont let this get to u. there is always time to ask ur teacher. dont take it too hard now

OpenStudy (anonymous):

Thanks. Night Habibi